OpenStudy (anonymous):

What is the exact value of cos 4/3 pi

OpenStudy (anonymous):

no calculator?

OpenStudy (anonymous):

-0.5

OpenStudy (anonymous):

look at the x-value at \[\frac{ 4\pi }{ 3 }\]

OpenStudy (anonymous):

0.999 or 1

OpenStudy (anonymous):

You mean in Rad or Deg? Rad : \[\frac{4}{3}\times\pi=\pi+\frac{1}{2}\times\frac{2}{3} \times\pi\] means \[cos((\pi)+(\frac{\pi}{3}))=cos(\pi)cos(\frac{\pi}{3})-sin(\pi)sin(\frac{\pi}{3})\] = \[(-1\times0.5)-(0\times \frac{\sqrt{3}}{2}) \]

OpenStudy (anonymous):

This is the first time I have used this site. I'm in my late 70's and foolishly said I'd help grandson with his maths. This is what I wanted to know. Thanks very much.

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

@HeilanGrannie it's my pleasure :)